# Entropy-conservative high-order methods for high-enthalpy gas flows

### v1

@@ 1 Introduction
Accurate simulation of non-equilibrium flows with chemical reactions and internal degrees of freedom is crucial for a multitude of aerospace applications [4, 32, 21]. For such flows, using higher-order discontinuous Galerkin (DG) methods is an attractive prospect, as it enables a more accurate simulation of turbulent effects, lessens the requirements on grid-shock alignment, improves computational efficiency due to data locality, and adds additional solution adaptivity capabilities via $p$-adaptivity [45, 23, 2, 31, 17, 34]. DG methods are also the natural framework for polynomial expansion-based uncertainty quantification and sensitivity analysis studies [43, 24], an area of research with growing importance, as larger and larger parts of the R&D cycle in aerospace are carried out via simulation tools. The less stringent requirements of DG methods with regards to shock-grid alignment [8] are also useful for uncertainty quantification and sensitivity analysis studies, as those oftentimes require automated simulation of hundreds and thousands of test cases, where automated well-aligned mesh generation is possible only for simple geometries [9].
@@
Development of DG methods suitable for aerospace applications is still an active area of research, as questions of numerical stability, shock capturing, and insufficient numerical diffusion inherent to higher-order methods need to be addressed. Recent achievements in these areas include the DG spectral element method with Legendre-Gauss-Lobatto quadratures (DGSEM-LGL) [11, 12, 39], artificial viscosity approaches [30, 8, 1], shock filtering [30, 3], and subcell limiting techniques [16, 26].
@@
Another question is the choice of the numerical flux functions: from a physics point of view, a crucial desired property of a numerical scheme is entropy conservation [44, 19, 18], especially in the context of DG methods [6, 36, 12]. For summation-by-part (SBP) schemes, this requires entropy-conservative numerical flux functions for the volume and surface fluxes [36, 12]. Augmented by a dissipative surface flux function, the spatial discretization then guarantees entropy stability [6]. Multiple entropy-conservative flux functions have been developed [36], but they usually assume restrictions on the description of the internal energy of the gas flow. Quite often, a constant ratio of specific heats is assumed. Whilst this greatly simplifies the derivation of numerical fluxes, it is not applicable to the kind of high-temperature flows encountered in real-life aerospace applications. In [20], the DG method was applied to the simulation of reacting mixtures with special attention paid to the numerical evaluation of fluxes that avoid generating spurious pressure oscillations. In a recent paper [34], an approach has been proposed to construct entropy stable flux functions for mixtures of molecules with vibrational spectra described by infinite harmonic oscillators. This significantly expands the range of applicability of such entropy-conservative schemes. However, the assumed infinite harmonic oscillator model is not the most accurate description of real-life vibrational spectra of molecules. In addition, the approach used to derive the flux functions does not easily generalize to other expressions for internal energy spectra.
@@
In the present work, a new approach to derive of entropy-conservative flux schemes is developed, based on interpolation of the specific heats and entropy integrals. It is applied to the simulation of inviscid high-enthalpy flows over a cylinder, and compared to results obtained with the second order finite volume-based DLR TAU solver [28].
@@ 2.1 Internal energy models
We now consider several models for $\varepsilon_{\mathrm{int}}(T)$. The simplest case is that of the calorically perfect gas, where $\varepsilon_{\mathrm{int}}(T)=c_{v}T$, where $c_{v}$ is the constant-valued specific heat at constant volume. The ratio of specific heats $\gamma=(c_{v}+k/m)/c_{v}$ in that case is also temperature-independent. The calorically perfect gas case, especially with $c_{v}=5/2k/m$ (corresponding $\gamma=1.4$), is the standard one in a large amount of literature concerned with CFD methods in general and entropy-conservative formulations in particular, and we do not consider it in detail.
In the present work, we assume that our gas consists of diatomic molecules, which possess only rotational and vibrational degrees of freedom. That is, we neglect the impact of electronic excitation: $\varepsilon_{\mathrm{int}}=\varepsilon_{\mathrm{tr}}+\varepsilon_{\mathrm{rot}}+\varepsilon_{\mathrm{vibr}}$, where $\varepsilon_{\mathrm{tr}}=\frac{3}{2}kT/m$ is the specific translational energy, and $\varepsilon_{\mathrm{rot}}$ and $\varepsilon_{\mathrm{vibr}}$ are the specific rotational and vibrational energies, respectively. Moreover, we assume a “rigid rotor” model [32]), that is, that the rotational and vibrational degrees of freedom are uncoupled. We also assume that the rotational mode is fully excited, leading to $\varepsilon_{\mathrm{rot}}=kT/m$. As a consequence, the specific heat of the rotational degrees of freedom simplifies to $c_{v,\mathrm{rot}}=k/m$. For a more detailed discussion of the validity of these assumptions, at least with regard to viscous fluid properties, the reader is referred to [25].
@@
[…]
@@ 2.2 Implementation of internal energy models in a CFD solver
Accounting for internal energy effects in a flow solver involves implementing several functions: 1) one to compute the flow temperature $T$ from the conservative variables, 2) one to compute a specific internal energy given the flow temperature $T$, 3) one to compute the specific heats given the flow temperature $T$, and 4) one to compute the flow entropy given the temperature and conservative variables. For multi-species flows and flows with thermal non-equilibrium (multi-temperature flows), these functions have to be extended accordingly to account for all the relevant dependencies on the partial densities and the temperatures of the internal degrees of freedom.
In general, no closed-form expression exists to compute $T$ from $\varepsilon_{\mathrm{int}}$. As such, existing packages for high-temperature flow simulation, such as Mutation++ [41], use a non-linear solver to obtain the flow temperature. In the present work, a single-species gas is considered, and an approach based on linear interpolation between tabulated values is used due to the lower computational cost; a discussion of the cost of using a non-linear solver in lieu of the linear interpolation is presented below.
First, reasonable minimum and maximum values of $T$ are chosen, e.g., $T_{\mathrm{min}}=10$ K, $T_{\mathrm{max}}=50000$ K, as well as a temperature discretization step $\Delta T$. The full specific internal energy $\varepsilon_{\mathrm{int}}=\varepsilon_{\mathrm{int}}(T)$ is computed over the specified temperature range with the corresponding step in $\Delta T$ and stored, as well as the total specific heat $c_{v}=c_{v}(T)$. Since one needs a way to obtain the temperature, a primitive variable, from energy, a conservative variable, the minimum and maximum internal energy values are computed as $\varepsilon_{\mathrm{int}}=\varepsilon_{\mathrm{int}}(T_{\mathrm{min}})$, $\varepsilon_{\mathrm{int}}=\varepsilon_{\mathrm{int}}(T_{\mathrm{max}})$, and the number of discretization points is chosen to be the same as for the discretization of the temperature range. Based on this, a uniform energy discretization step $\Delta\varepsilon$ is computed, and for each energy in the range, the corresponding temperature is found using a non-linear solver. As this step is done only once during the initialization of the flow solver, a high accuracy can be used for the non-linear solve, as even for a large number of discretization points, the cost taken up by the pre-processing is negligible compared to the time taken up by the flow simulation itself. With these tabulated arrays, reconstructing temperature from the energy and computing energy and/or specific heat from the temperature requires only finding the closest position in the array. This is achieved via division by the discretization step and subsequent rounding down to an integer, and performing a linear interpolation between the nearest values in the array. In case of a multi-species flow, where the inference of temperature from energy requires solving a non-linear equation, the linear interpolation of energy and specific heats is still utilized in the non-linear solver to avoid a potentially expensive computation of energy and its derivative with respect to the temperature at each iteration.
@@ 3.1 Entropy and entropy variables
$s=\int_{0}^{T}\frac{c_{v}(\tau)}{\tau}\mathrm{d}\tau-\frac{k}{m}\ln\rho=\eta(T)-\frac{k}{m}.$ (7)
@@
Using the chain rule [13], we can write the entropy variable vector
@@ 3.2 Derivation of entropy-conservative flux
where $\llbracket a\rrbracket$ = $a_{+}-a_{-}$ is the jump in the value of a flow variable $a$, and $a_{-}$ and $a_{+}$ denote the values of a flow variable $a$ at two points between which the flux is computed. The flux potentials are given by [34]:
@@
To obtain the flux $\mathbf{F}^{\mathrm{num},EC,j}$, we need to write $\llbracket\omega\rrbracket$ and $\llbracket\psi_{j}\rrbracket$ in terms of jumps in the primitive variables
@@
$\mathbf{F}^{\mathrm{num},EC,j}=\left(A^{-1}\left(\mathbf{z}_{-},\mathbf{z}_{+}\right)B\left(\mathbf{z}_{-},\mathbf{z}_{+}\right)\right)^{\mathrm{T}}.$ (22)
In order to obtain rewrite the jumps in the entropy variables in terms of jumps in the primitive variables, we first need to introduce the following averaging operators [36]
@@
Then, the following relations hold [36]:
@@
All that is left is to express the jumps in $\eta(T)$ and $\varepsilon_{\mathrm{int}}(T)$ in terms of jumps in $T$. At first this seems impossible unless some specific expressions for $c_{v}(T)$ and $\varepsilon_{\mathrm{int}}$ are provided. Even then, a closed-form solution is not necessarily available. For the simplest case of an infinite harmonic oscillator [34], the resulting flux formulae are quite complex and involve hyperbolic functions. Another approach is to use the NASA polynomials [29], which approximate specific heats as polynomials in $T$ and $1/T$. The polynomial fits for entropies and enthalpies are obtained via simple analytical integration. One can plug in the polynomial approximations into (32) and obtain a closed-form expression for the jump in $\omega_{1}$ using the relevant chain rules, which is the procedure employed in [13]. The drawback of the approach is that for each internal energy model, a new set of fits has to be performed as the coefficients given by NASA take into account effects such as anharmonicity, electronic excitatation, etc. Moreover, in thermally non-equilibrium multi-temperature flows, additional fits for the vibrational components of the internal energies need to be provided.
@@
Here $T^{\ast}$, $T^{\ast\ast}$ are some values of the temperature in the interval $\left[\mathrm{min}\left(T_{-},T_{+}\right),\mathrm{max}\left(T_{-},T_{+}\right)\right]$. But how do we find the values ${c_{v}\left(T^{\ast}\right)}/{T^{\ast}}$, $c_{v}\left(T^{\ast\ast}\right)$?
The following algorithm is proposed: given $\mathbf{u}_{+}$, $\mathbf{u}_{-}$, we first compute $\left\llbracket\varepsilon_{\mathrm{int}}\right\rrbracket$ and $\left\llbracket\eta(T)\right\rrbracket$. The jump in the internal energy can be computed explicitly. To compute the jump in the integral part of the entropy $\eta(T)$, we pre-compute tabulated values of $\eta(T)$ with a step size of $\Delta T$ and store them. Thus, once we convert the conservative variables to the primitive variables, we obtain $T_{+}$ and $T_{-}$, and use them to find via linear interpolation the jump in $\eta(T)$. Since we already know $\llbracket T\rrbracket$, we simply divide the left-hand sides of (33), (34) by $\llbracket T\rrbracket$ and obtain the required values.
Thus, we can write out the final expression for $\left\llbracket\omega_{1}\right\rrbracket$:
@@
3.3 Limiting case
We can check that in the case of $c_{v}=\mathrm{const}$, the formulas reduce to the fluxes as given by expression (75) in [36]. For the density and velocity fluxes, this is immediately evident if one uses $R_{\mathrm{specific}}T$ instead of $T$ as the variable, where $R_{\mathrm{specific}}=\frac{k}{m}$. All that remains is to simplify the expressions for the energy flux (we only consider the flux in the $y$ direction, the derivation for the $x$ direction is done by analogy). For constant $c_{v}$, we have that $\eta(T_{+})-\eta(T_{-})=\int_{T_{-}}^{T_{+}}{c_{v}}/{\tau}\mathrm{d}\tau=c_{v}\left(\log T_{+}-\log T_{-}\right)$. We can therefore write that $T^{\ast}=\left\{\!\left\{T\right\}\!\right\}_{\mathrm{log}}$. We also have that $\left\{\!\left\{\varepsilon_{\mathrm{int}}\right\}\!\right\}=c_{v}\left\{\!\left\{T\right\}\!\right\}$. Therefore, we can re-write expression (43) as

### v2

@@ 1 Introduction
Accurate simulation of non-equilibrium flows with chemical reactions and internal degrees of freedom is crucial for a multitude of aerospace applications [6, 40, 29]. For such flows, using higher-order discontinuous Galerkin (DG) methods is an attractive prospect, as it enables a more accurate simulation of turbulent effects, lessens the requirements on grid-shock alignment, improves computational efficiency due to data locality, and adds additional solution adaptivity capabilities via $p$-adaptivity [54, 23, 31, 4, 39, 25, 42]. DG methods are also the natural framework for polynomial expansion-based uncertainty quantification and sensitivity analysis studies [52, 32], an area of research with growing importance, as larger and larger parts of the R&D cycle in aerospace are carried out via simulation tools. The less stringent requirements of DG methods with regards to shock-grid alignment [11] are also useful for uncertainty quantification and sensitivity analysis studies, as those oftentimes require automated simulation of hundreds and thousands of test cases, where automated well-aligned mesh generation is possible only for simple geometries [12].
@@
Development of DG methods suitable for aerospace applications is still an active area of research, as questions of numerical stability, shock capturing, and insufficient numerical diffusion inherent to higher-order methods need to be addressed. Recent achievements in these areas include the DG spectral element method with Legendre-Gauss-Lobatto quadratures (DGSEM-LGL) [16, 17, 46], artificial viscosity approaches [38, 11, 3], shock filtering [38, 5], and subcell limiting techniques [24, 48, 34, 47].
@@
Another question is the choice of the numerical flux functions: from a physics point of view, a crucial desired property of a numerical scheme is entropy conservation [53, 27, 26, 14], especially in the context of DG methods [9, 8, 44, 17]. For summation-by-part (SBP) schemes, ensuring entropy conservation/entropy stability at the semi-discrete level requires entropy-conservative numerical flux functions for the volume and surface fluxes [44, 17]. Augmented by a dissipative surface flux function, the spatial semi-discretization then guarantees entropy stability [8]. Multiple entropy-conservative flux functions have been developed [44], but they usually assume restrictions on the description of the internal energy of the gas flow. Quite often, a constant ratio of specific heats is assumed. Whilst this greatly simplifies the derivation of numerical fluxes, it is not applicable to the kind of high-temperature flows encountered in real-life combustion and aerospace applications, where excitation internal degrees of freedom plays a significant role [6, 37, 40, 29]. Therefore, development of new numerical fluxes that possess attractive properties, such as preservation of entropy, kinetic energy, and pressure equilibrium, for gases with non-trivial internal energy functions and/or equations of state is of very high relevance to the applied CFD community in order to drive the development of new robust higher-order methods for large-scale simulations.
Several works have concerned themselves with exactly such developments. In [19, 10], entropy-conservative schemes were developed for gases with internal degrees of freedom were the internal energy is modelled by a polynomial function of temperature [37]. In [28], the DG method was applied to the simulation of reacting mixtures with special attention paid to the numerical evaluation of fluxes that avoid generating spurious pressure oscillations; the internal energies were also modelled using NASA polynomials. In [2], entropy-conservative discretizations have been proposed for gases with an arbitrary (non-ideal gas law) equation of state; the impact of internal degrees of freedom on the ideal gas effects were also modelled by polynomial functions of temperature. In [21], entropy-stable flux functions were derived for gas mixtures with internal energies described either by polynomial fits, or with vibrational spectra modelled by the infinite harmonic oscillator model. Flux functions for mixtures with thermal non-equilibrium between the vibrational and rotational-translational modes were derived as well. Similarly, in [42], an approach has been proposed to construct entropy stable flux functions for mixtures of molecules with vibrational spectra described by infinite harmonic oscillators. In [1] an alternative approach for evaluating entropy-conservative fluxes for gases with arbitrary equations of state was proposed, based on computation of correction terms as solutions of an optimization problem.
Thus, recent advancements in the field of structure-preserving methods for gas dynamics have significantly expanded the range of applicability of such entropy-conservative schemes to real gases with complex equations of state and dependencies of internal energy on the gas temperature. However, they usually still rely either on polynomial fits or the infinite harmonic oscillator model, which is not the most accurate description of real-life vibrational spectra of molecules. In addition, the approaches used to derive the flux functions does not easily generalize to other expressions for internal energy spectra.
@@
In the present work, a new approach to derive entropy-conservative flux schemes is developed, based on interpolation of the specific heats and entropy integrals. It is applied to the simulation of inviscid high-enthalpy flows over a cylinder, and compared to results obtained with the second order finite volume-based DLR TAU solver [36].
@@ 2.1 Internal energy models
We now consider several models for $\varepsilon_{\mathrm{int}}(T)$. The simplest case is that of the calorically perfect gas, where $\varepsilon_{\mathrm{int}}(T)=c_{v}T$, where $c_{v}$ is the constant-valued specific heat at constant volume. The ratio of specific heats $\gamma=(c_{v}+k/m)/c_{v}$ in this case is also temperature-independent. The calorically perfect gas case, especially with $c_{v}=5/2k/m$ (corresponding to $\gamma=1.4$), is the standard one in a large amount of literature concerned with CFD methods in general and entropy-conservative formulations in particular, and we do not consider it in detail. The factor $5/2$ comes of the contribution of the translational degrees of freedom ($3/2$) and the fully excited rotational degrees of freedom ($/2$ 2) for molecular gases.
In the present work, we assume that our gas consists of diatomic molecules, which possess only rotational and vibrational degrees of freedom. That is, we neglect the impact of electronic excitation and can write $\varepsilon_{\mathrm{int}}=\varepsilon_{\mathrm{tr}}+\varepsilon_{\mathrm{rot}}+\varepsilon_{\mathrm{vibr}}$, where $\varepsilon_{\mathrm{tr}}=\frac{3}{2}kT/m$ is the specific translational energy, and $\varepsilon_{\mathrm{rot}}$ and $\varepsilon_{\mathrm{vibr}}$ are the specific rotational and vibrational energies, respectively. Moreover, we assume a “rigid rotor” model [40], that is, that the rotational and vibrational degrees of freedom are uncoupled. We also assume that the rotational mode is fully excited, leading to $\varepsilon_{\mathrm{rot}}=kT/m$. As a consequence, the specific heat of the rotational degrees of freedom simplifies to $c_{v,\mathrm{rot}}=k/m$. For a more detailed discussion of the validity of these assumptions, at least with regard to viscous fluid properties, the reader is referred to [33].
@@
Figure 1 shows the impact the choice of the vibrational energy spectrum description has on the thermodynamic properties of the gas (in this case, the heat capacity ratio $\gamma$). Three different models of increasing complexity are considered: the infinite harmonic oscillator, the cut-off harmonic oscillator, and the cut-off anharmonic oscillator. For both of the cut-off models, the cut-off was determined by the dissociation energy $D$. It can be seen that the heat capacity ratio strongly deviates from the room temperature value of 1.4, with all the cut-off models exhibiting non-monotone behaviour. For nitrogen, the effects of vibrational excitation and vibrational spectrum cut-off appear at higher temperatures than for oxygen, due to its higher characteristic vibrational temperature and larger dissociation energy. Due to the larger dissociation energy of nitrogen, the difference between the infinite and cut-off harmonic oscillator models is also not significant in the temperature range considered, appearing only for $T>12000$ K, whereas for oxygen the difference becomes noticeable already for temperatures above 6000 K. The impact of anharmonicity is more pronounced for both molecules. It can thus be concluded that whilst the infinite harmonic oscillator model does allow for simulation of more complex thermodynamical phenomena compared to a calorically perfect gas model, it is noticeably less accurate than the more detailed cut-off oscillator models for high temperatures. Note that the presented results do not account for effects such as electronic excitation, which further complicates the thermodynamic behaviour of gases at temperatures higher than 7000-10000 K [37, 7].
@@ 2.2 Implementation of internal energy models in a CFD solver
Accounting for internal energy effects in a flow solver involves implementing several functions: 1) one to compute the flow temperature $T$ from the conservative variables, 2) one to compute a specific internal energy given the flow temperature $T$, 3) one to compute the specific heats given the flow temperature $T$, and 4) one to compute the flow entropy given the temperature and conservative variables.
In general, no closed-form expression exists to compute $T$ from $\varepsilon_{\mathrm{int}}$. As such, existing packages for high-temperature flow simulation, such as Mutation++ [50], use a non-linear solver to obtain the flow temperature. In the present work, a single-species gas is considered, and an approach based on linear interpolation between tabulated values is used due to the lower computational cost.
First, reasonable minimum and maximum values of $T$ are chosen, e.g., $T_{\mathrm{min}}=10$ K, $T_{\mathrm{max}}=50000$ K, as well as a temperature discretization step $\Delta T$. The full specific internal energy $\varepsilon_{\mathrm{int}}=\varepsilon_{\mathrm{int}}(T)$ is computed over the specified temperature range with the corresponding step in $\Delta T$ and stored, as well as the total specific heat $c_{v}=c_{v}(T)$. This allows for efficient computation of temperature, a primitive variable, from energy, a conservative variable, as the Newton solver used to solve for temperature can use the linear interpolation for $c_{v}$ and $\varepsilon_{\mathrm{int}}$ instead of the potentially computationally more expensive analytical expressions.
@@ 3.1 Entropy and entropy variables
$s=\int_{0}^{T}\frac{c_{v}(\tau)}{\tau}\mathrm{d}\tau-\frac{k}{m}\ln\rho=\eta(T)-\frac{k}{m}\ln\rho.$ (7)
@@
Using the chain rule, we can write the entropy variable vector [22]
@@ 3.2 Derivation of entropy-conservative flux
where $\llbracket a\rrbracket$ = $a_{+}-a_{-}$ is the jump in the value of a flow variable $a$, and $a_{-}$ and $a_{+}$ denote the values of a flow variable $a$ at two points between which the flux is computed, that is, the flux $\mathbf{F}^{\mathrm{num},EC,j}$ depends on both flow states $\mathbf{u}_{\pm}$. The flux potentials are given by [42]:
@@
To obtain the flux $\mathbf{F}^{\mathrm{num},EC,j}$, we need to write $\llbracket\omega\rrbracket$ and $\llbracket\psi_{j}\rrbracket$ in terms of jumps our preferred variables (in this case, the primitive variables $\mathbf{z}$)
@@
$\mathbf{F}^{\mathrm{num},EC,j}=\left(B\left(\mathbf{z}_{-},\mathbf{z}_{+}\right)A^{-1}\left(\mathbf{z}_{-},\mathbf{z}_{+}\right)\right)^{\mathrm{T}}.$ (22)
In order to obtain rewrite the jumps in the entropy variables in terms of jumps in the primitive variables, we first need to introduce the following averaging operators [26, 44]
@@
Then, the following relations hold:
@@
All that is left is to express the jumps in $\eta(T)$ and $\varepsilon_{\mathrm{int}}(T)$ in terms of jumps in $T$. At first this seems impossible unless some specific expressions for $c_{v}(T)$ and $\varepsilon_{\mathrm{int}}$ are provided. Even then, a closed-form solution is not necessarily available. For the simplest case of an infinite harmonic oscillator [42], the resulting flux formulae are quite complex and involve hyperbolic functions. Another approach is to use the NASA polynomials [37], which approximate specific heats as polynomials in $T$ and $1/T$. The polynomial fits for entropies and enthalpies are obtained via simple analytical integration. One can plug in the polynomial approximations into (32) and obtain a closed-form expression for the jump in $\omega_{1}$ using the relevant chain rules, which is the procedure employed in [19]. The drawback of the approach is that for each internal energy model, a new set of fits has to be performed as the coefficients given by NASA take into account effects such as anharmonicity, electronic excitation, etc. Moreover, in thermally non-equilibrium multi-temperature flows, additional fits for the vibrational components of the internal energies need to be provided.
@@
Here $T^{\ast}$, $T^{\ast\ast}$ are some values of the temperature in the interval $\left[\mathrm{min}\left(T_{-},T_{+}\right),\mathrm{max}\left(T_{-},T_{+}\right)\right]$. Based on this formalism, we can simply define the following quantities:
$\frac{c_{v}\left(T^{\ast}\right)}{T^{\ast}}\coloneq\frac{\left\llbracket\eta(T)\right\rrbracket}{\llbracket T\rrbracket},$ (35)
$c_{v}\left(T^{\ast\ast}\right)\coloneq\frac{\left\llbracket\varepsilon_{\mathrm{int}}\right\rrbracket}{\llbracket T\rrbracket}.$ (36)
With that, we can write out the final expression for $\left\llbracket\omega_{1}\right\rrbracket$:
@@
The only remaining part is the computation of $\frac{c_{v}\left(T^{\ast}\right)}{T^{\ast}}$, $c_{v}\left(T^{\ast\ast}\right)$. It should be stressed that up to this point, no quantities have been approximated. For practical purposes, however, one requires some numerical procedures to compute $\llbracket T\rrbracket$ and $\left\llbracket\eta(T)\right\rrbracket$. For example, given an analytical expression for $\varepsilon_{\mathrm{int}}(T)$ (e.g. one incorporating the infinite harmonic oscillator model (5) or cut-off oscillator model (2) for the internal energy), one can use Newton’s method to compute $T_{-}$, $T_{+}$ from the given values of $\varepsilon_{\mathrm{int},-}$, $\varepsilon_{\mathrm{int},+}$ to a desired degree of precision. From the computed temperature values, one can then evaluate $\eta(T_{-})$ and $\eta(T_{+})$ numerically using some quadrature rule. The disadvantage of such an approach is the potentially high computational cost, as each flux evaluation requires multiple evaluations of $\varepsilon_{\mathrm{int}}(T)$ and $c_{v}(T)$ in the Newton solver iterations and the numerical integration for $\eta(T)$.
In the present work, we propose the following procedure. Values of $\varepsilon_{\mathrm{int}}$ and $c_{v}(T)$ are pre-computed and tabulated at the start of a simulation over a range of temperatures with a uniform step size of $\Delta T$. During the course of a simulation, piece-wise linear interpolation is then used to compute the values of $\varepsilon_{\mathrm{int}}$, $c_{v}(T)$ where required. The piece-wise linear interpolation is of $c_{v}(T)$ is used to compute $\eta(T)$ exactly. We can write
$\eta(T)=\sum_{i=0}^{N-1}\eta_{i}+\left(c_{v,N}-\frac{(c_{v}(T)-c_{v,N})T_{N}}{\Delta T}\right)\ln\left(\frac{T}{T_{N}}\right)+\left(c_{v}(T)-c_{v,N}\right).$ (46)
Here $c_{v,i}$ denotes a tabulated value of $c_{v}(T)$ computed at a tabulated temperature of $T_{i}$, $c_{v}(T)$ is a linearly interpolated value of $c_{v}$ at an arbitrary temperature. $N$ is defined as $\lfloor T-T_{\mathrm{min}}/\Delta T\rfloor$, where $T_{min}$ is the minimum temperature used for pre-computing the values. The values $\eta_{i}$ are the integrals of $c_{v}(T)/T$ computed over the $\Delta T$-sized intervals assuming a linear interpolation of $c_{v}(T)$ over the interval:
$\eta_{i}=\left(c_{v,i}-\frac{(c_{v,i+1}-c_{v,i})T_{i}}{\Delta T}\right)\ln\left(\frac{T_{i+1}}{T_{i}}\right)+\left(c_{v,i+1}-c_{v,i}\right).$ (47)
To reduce the computational cost of computing $\eta(T)$ during a simulation, the values of $\sum_{i=0}^{N}\eta_{i}$ can also be tabulated for $N=0,\ldots N_{\mathrm{max}}$, and the computation of $\eta(T)$ is then reduced to getting a value from a look-up table and evaluating one logarithm. The proposed algorithm not only provides a direct numerical procedure for evaluation of the entropy-conservative flux presented above for a gas with an arbitrary equation for the internal energy distribution function, but also potentially leads to a reduction in the computational cost (as potentially expensive on-the-fly computations are replaced by linear interpolation). In principle, other higher-order interpolation methods can be used, as the developed flux formulation is independent of the numerical procedures used to estimate $\frac{c_{v}\left(T^{\ast}\right)}{T^{\ast}}$, $c_{v}\left(T^{\ast\ast}\right)$. The error due to the interpolation of energy and specific heats will be analyzed in the next section, and in the numerical results section it will be shown that with a reasonable choice of $\Delta T$ the error in the entropy production rate can be made to be on the order of machine precision, thus guaranteeing numerical entropy conservation.
@@
3.3 Properties of the entropy-conservative flux
We now investigate and prove several properties of the flux (38)–(45).
Theorem 1.
The numerical flux given by Eqns. (38)–(45) is consistent if $\lim_{\begin{subarray}{c}T_{-}\to T\\ T_{+}\to T\end{subarray}}{c_{v}\left(T^{\ast}\right)}/{T^{\ast}}={c_{v}(T)}/{T}$ and $\lim_{\begin{subarray}{c}T_{-}\to T\\ T_{+}\to T\end{subarray}}c_{v}\left(T^{\ast\ast}\right)=c_{v}(T)$.
Proof.
We prove consistency of the flux in the $x$ spatial direction, as consistency of the flux in the $y$ direction can be proved by analogy. We need to show that $\mathbf{F}^{\mathrm{num},x}\left(\mathbf{u},\mathbf{u}\right)=\mathbf{f}_{x}(\mathbf{u})$.
The averaging operators (23)–(25) are consistent, thus immediately leading to consistency of the density and momentum fluxes ${F}^{\mathrm{num},x}_{\rho}$, ${F}^{\mathrm{num},x}_{\rho v_{x}}$, ${F}^{\mathrm{num},x}_{\rho v_{y}}$.
We now consider the energy flux in more detail. Per the requirement that ${c_{v}\left(T^{\ast}\right)}/{T^{\ast}}\to{c_{v}(T)}/{T}$ and $c_{v}\left(T^{\ast\ast}\right)\to c_{v}(T)$ as $T_{+}\to T$ and $T_{-}\to T$ we have that ${c_{v}\left(T^{\ast}\right)}/{T^{\ast}}-\left\{\!\left\{1/T\right\}\!\right\}c_{v}\left(T^{\ast\ast}\right)=0$.
Using the consistency of the averaging operators we can therefore write
$F^{\mathrm{num},x}_{E}=F^{\mathrm{num},x}_{\rho}\left(\varepsilon_{\mathrm{int}}-\frac{v_{x}^{2}+v_{y}^{2}}{2}\right)+v_{x}F^{\mathrm{num},x}_{\rho v_{x}}+v_{y}F^{\mathrm{num},x}_{\rho v_{y}}.$ (48)
Applying the consistency of the density and momentum fluxes, the energy flux can be rewritten as
$F^{\mathrm{num},x}_{E}=\rho v_{x}\left(\varepsilon_{\mathrm{int}}-\frac{v_{x}^{2}+v_{y}^{2}}{2}\right)+\rho v_{x}^{3}+v_{x}p+\rho v_{x}v_{y}^{2}=v_{x}\left(\rho\varepsilon_{\mathrm{int}}+\rho\frac{v_{x}^{2}+v_{y}^{2}}{2}+p\right)=v_{x}\left(E+p\right).$ (49)
This can be seen to be exactly equal to the energy flux in the $x$ direction for Euler equations as given by (3). ∎
Theorem 2.
The numerical flux given by Eqns. (38)–(45) is kinetic energy preserving.
Proof.
We consider the flux in the $x$ direction, proof for the $y$ direction is analogous. The momentum flux for $x$ component of the momentum is given by
$F^{\mathrm{num},x}_{\rho v_{x}}=F^{\mathrm{num},x}_{\rho}\left\{\!\left\{v_{x}\right\}\!\right\}+\frac{k}{m}\frac{\left\{\!\left\{\rho\right\}\!\right\}}{\left\{\!\left\{1/T\right\}\!\right\}}.$ (50)
The quantity $\frac{k}{m}\frac{\left\{\!\left\{\rho\right\}\!\right\}}{\left\{\!\left\{1/T\right\}\!\right\}}$ is a consistent approximation of the pressure, and thus, the numerical flux (38)–(45) is kinetic energy preserving [27, 9, 44]. ∎
Remark 1.
In the case of $c_{v}=\mathrm{const}$, the flux given by Eqns. (38)–(45) and evaluated using the proposed procedure (piece-wise linear interpolation of $c_{v}(T)$ and subsequent use of Eqn. 46) reduces to the flux proposed by Chandrashekar [9], differing only up to a trivial change of variables from $T$ to $R_{\mathrm{specific}}T$.
Proof.
For the density and velocity fluxes, this is immediately evident if one uses $R_{\mathrm{specific}}T$ instead of $T$ as the variable, where $R_{\mathrm{specific}}=\frac{k}{m}$. All that remains is to simplify the expressions for the energy flux (we only consider the flux in the $y$ direction, the derivation for the $x$ direction is done by analogy). For constant $c_{v}$, we have that $\eta(T_{+})-\eta(T_{-})=\int_{T_{-}}^{T_{+}}{c_{v}}/{\tau}\mathrm{d}\tau=c_{v}\left(\log T_{+}-\log T_{-}\right)$. Since this is a specific case of a piece-wise linear $c_{v}$, this analytical integration is recovered exactly in our proposed numerical procedure. We can therefore write that $T^{\ast}=\left\{\!\left\{T\right\}\!\right\}_{\mathrm{log}}$. We also have that $\left\{\!\left\{\varepsilon_{\mathrm{int}}\right\}\!\right\}=c_{v}\left\{\!\left\{T\right\}\!\right\}$. Therefore, we can re-write expression (45) as